# Investigating LLMs in Clinical Triage: Promising Capabilities, Persistent Intersectional Biases

### v1

@@ Abstract
Large Language Models (LLMs) have shown promise in clinical decision support, yet their application to triage remains underexplored. We systematically investigate the capabilities of LLMs in emergency department triage through two key dimensions: (1) robustness to distribution shifts and missing data, and (2) counterfactual analysis of intersectional biases across sex and race. We assess multiple LLM-based approaches, ranging from continued pre-training to in-context learning, as well as machine learning approaches. Our results indicate that LLMs exhibit superior robustness, and we investigate the key factors contributing to the promising LLM-based approaches. Furthermore, in this setting, we identify gaps in LLM preferences that emerge in particular intersections of sex and race. LLMs generally exhibit sex-based differences, but they are most pronounced in certain racial groups. These findings suggest that LLMs encode demographic preferences that may emerge in specific clinical contexts or particular combinations of characteristics.
@@ 1 Introduction
Large language models (LLMs) have shown promise for various clinical settings (Cascella et al. 2023; Park et al. 2024; Li et al. 2024c; Jiang et al. 2023), applications ranging from clinical decision support (Liu et al. 2023) to conversational systems (Montagna et al. 2023). Amongst these applications, emergency departments (EDs) are among the most dynamic and high-stakes clinical environments, serving as the initial point of contact for patients requiring urgent medical attention (Franc et al. 2024). A critical component of ED operations is the triage process, where patients are prioritized based on clinical urgency to ensure timely and appropriate care. Overtriage misallocates scarce resources in a time of department overcrowding and undertriage endangers time-sensitive patients (Chmielewski and Moretz 2022). In an era of historic ED overcrowding, these issues are a growing concern. While triage is an imperfect practice due to limited information (Hinson et al. 2018), triage practices performed by humans are further prone to challenges such as variability in decision-making, lack of training programs, cognitive biases, and decreased quality due to overwhelming patient arrivals (Joseph et al. 2023; Chmielewski and Moretz 2022).
@@
Traditional machine learning (ML) algorithms have performed well (Hong, Haimovich, and Taylor 2018; Goto et al. 2019; Miles et al. 2020), but they are limited by their reliance on structured input data and sensitivity to distributional shifts. Health data in these settings predominantly consists of unstructured text, such as clinical notes, and significantly different triage distributions have been observed across hospitals (Chmielewski and Moretz 2022) which can present major challenges to ML-based methods (Li et al. 2024c) . In contrast, the clinical knowledge (Singhal et al. 2023) of LLMs and their reasoning abilities pose them as a robust model, able to adapt to diverse settings. Moreover, LLMs can provide explanations that act as a second opinion for practitioners, mitigating potential mispractice (Goh et al. 2023). LLMs offer a promising approach to make triage in the ED more efficient, accurate, and equitable (Preiksaitis et al. 2024). Hence, the demand for artificial intelligence (AI) assistance in this process is substantial (Friedman, Delgado, and Weissman 2024), and it is already being deployed  (Johns Hopkins Medicine 2022).
Despite their potential, the application of LLMs to triage and their robustness remains under-studied. Only a few recent studies have examined their capabilities (Williams et al. 2024b; Masanneck et al. 2024; Lu et al. 2024). A significant opportunity lies in exploring LLM-based approaches to improve both prediction and clinical utility. Existing studies have predominantly focused on basic prompting techniques without testing advanced methodologies. In addition, we find that most studies have not replicated realistic triage scenarios, as noted by (Franc et al. 2024), as it is usually limited to a brief description of their symptoms, pain, and vital signs. There are even fewer studies, if any exist, on AI in alternative triage settings, such as telephone triage which involves assessing a patient’s health concerns over the phone.
Furthermore, a concern comes from the potential biases and inconsistencies in LLMs when applied to triage decisions, where unfair prioritization leads to adverse outcomes against certain demographics. While research has documented various forms of bias in LLMs across different healthcare contexts (Guevara et al. 2024; Omiye et al. 2023; Zack et al. 2024; Yang et al. 2024), the implications for hospital triage remain unclear. This knowledge gap is critical given that triage can affect vulnerable patient populations. The high-stakes nature of triage decisions necessitates an extensive assessment of LLMs’ performance across diverse patient populations and clinical scenarios before any potential deployment in real-world settings.
@@
    We comprehensively evaluate across distribution shifts, missing data, and datasets ranging from real hospital records to curated case studies. We assess various LLM-based approaches, ranging from continued pre-training to in-context learning, providing insights to for future research directions.
@@ 2 Related Works
LLMs in Biomedical and Clinical Applications. LLMs have demonstrated impressive abilities across diverse scenarios in the biomedical and clinical domain, including fine-tuning general LLMs with biomedical-specific corpora (Yunxiang et al. 2023; Shu et al. 2023; Chen et al. 2023; Zhang et al. 2023; Labrak et al. 2024) and equipping frozen LLMs with techniques like chain-of-thought (CoT) reasoning and retrieval-augmented generation (RAG) (Jeong et al. 2024; Toufiq et al. 2023; Lee et al. 2024; Li et al. 2024a; Zakka et al. 2024; Lozano et al. 2023; Shang et al. 2024). In our study, we focus on a few of these recent approaches that have yet to be applied to triage. For example, studies have shown that careful selection of examples can significantly impact model performance (Rubin, Herzig, and Berant 2021) and few-shot prompting techniques, such as retrieving k-nearest neighbors (KNN), have been proposed to enhance LLM performance by retrieving relevant examples (Xu et al. 2023). Furthermore, recent studies have explored how LLMs acquire knowledge and expertise (Ovadia et al. 2023; Zhou et al. 2023) and we extend these investigations to our specific setting.
@@
LLMs in the Emergency Department. Recent studies have explored LLMs in triage tasks. For example, a study evaluated the performance of LLMs on a binary triage task, comparing two patients and determining which presented higher acuity. However, their approach relied on clinical notes generated after the fact  (Williams et al. 2024b). Another study assessed triage accuracy using CoT prompting and found that conditioning on demographic information did not affect model performance. However, the limited scope of their dataset, 45 vignettes, restricts the generalizability of their findings  (Ito et al. 2023). Williams extended the study of LLMs to emergency department scenarios such as predicting hospital admission, the need for radiological investigations, and antibiotic prescriptions. The models’ performance remained significantly lower than that of resident physicians across most tasks  (Williams et al. 2024a). The most advanced work to date is a retrieval-augmented LLM-based agent system that leverages the Emergency Severity Index (ESI) handbook  (Lu et al. 2024). However, their dataset, derived from the handbook, is inherently limited by its small sample size and idealized scenarios, and their code and results are currently un-replicatable. In real-world triage, information is limited and frequently incomplete, posing challenges for LLMs operating in dynamic clinical settings. We believe this calls for better benchmarking, which has recently emerged (Li et al. 2024d; Ouyang et al. 2024). These works represent a crucial step forward, but further research is needed to evaluate the readiness of LLMs.
@@
LLM Bias in Clinical Settings. Recent studies have highlighted significant biases in LLMs, particularly within clinical contexts. Hofmann et al. (2024) revealed that LLMs often exhibit covert rather than overt biases, making such biases less detectable and potentially more harmful. Their study identified biases against individuals speaking African American English dialects, illustrating the subtle ways in which LLMs can perpetuate inequities. Zack et al. (2024) found that when asking GPT-4 to describe a case of sarcoidosis, the model produced a vignette of a Black patient 97% of the time and a Black female patient 81% of the time. Similarly, Yang et al. (2024) employed counterfactual analysis to demonstrate that LLMs’ biased behaviors often align with existing healthcare disparities. For instance, white patients were systematically provided better treatment recommendations than Black patients, despite identical clinical information. However, to date, no research has systematically explored the intersectionality of sex and race in these biases—a critical gap that we aim to address in this work.
3 Experiments
@@ 3.1 Evaluating the Clinical Utility of LLMs
Datasets. We utilize three datasets: (1) hospital records from MIMIC-IV-ED v2.2 (Johnson et al. 2023), which follow the ESI protocol; (2) hospital records using the Korean Triage Acuity System (KTAS) (Moon et al. 2019) which were subsequently re-labeled in hindsight by three experts; and (3) curated case studies from the ESI handbook (Gilboy 2005). Notably, these datasets differ in the reliability and nature of their ground truth labels, which affects how we recommend interpreting results based on them.
@@
These datasets differ in two key dimensions. First, label reliability forms a gradient: MIMIC labels reflect nurses’ initial assessments, which are quite prone to mistakes; KTAS labels benefit from expert consensus that was formed after the fact; and the ESI handbook comprises of curated case studies with unambiguous acuity assignments based on established protocols. Second, information availability varies, with KTAS containing additional structured features beyond MIMIC, and the ESI handbook providing the most comprehensive patient profiles.
Benchmark Creation. We construct our benchmark with carefully controlled train-test distributions. For MIMIC-IV-ED, we partition data by temporal cohorts, using 10,000 patients from 2014–2016 for training and 1,000 patients from 2017–2019 for testing. Given the dataset’s scale, we apply stratified sampling across acuity levels and missingness to ensure representation of incomplete records. For KTAS, we introduce a domain shift by training on 689 patients from a local hospital and testing on 580 patients from a separate regional hospital—though both are academic, urban medical centers. Finally, for the ESI handbook, we adopt the split from Lu et al. (2024).
ESI-Handbook KTAS MIMIC
Methodology QWK MSE ()($\downarrow$)( ↓ ) F-1 Acc. QWK MSE ()($\downarrow$)( ↓ ) F-1 Acc. QWK MSE ()($\downarrow$)( ↓ ) F-1 Acc.
LogReg N.A. N.A. N.A. N.A. 0.369 0.813 0.418 0.472 0.306 0.550 0.483 0.564
XGBoost N.A. N.A. N.A. N.A. 0.059 1.475 0.291 0.342 0.216 1.616 0.211 0.210
BioBERT 0.117 2.904 0.251 0.287 0.074 1.169 0.261 0.330 0.498 0.432 0.640 0.656
Vanilla 0.923 0.213 0.788 0.787 0.599 0.634 0.511 0.520 0.562 0.561 0.585 0.568
AutoCoT 0.932 0.185 0.816 0.815 0.568 0.668 0.493 0.506 0.486 0.665 0.581 0.566
CoT 0.909 0.225 0.809 0.809 0.598 0.660 0.520 0.522 0.386 0.734 0.514 0.510
FewShot (40 shots) 0.940 0.169 0.826 0.831 0.603 0.639 0.516 0.518 0.546 0.494 0.567 0.563
FewShotCoT (40 shots) 0.894 0.292 0.708 0.708 0.624 0.573 0.559 0.561 0.550 0.486 0.591 0.588
KATE (40 shots) 0.934 0.180 0.821 0.820 0.667 0.413 0.696 0.718 0.654 0.352 0.677 0.678
KATECoT (40 shots) 0.917 0.236 0.763 0.764 0.667 0.497 0.607 0.608 0.601 0.402 0.639 0.640
Llama-8B (Fine-tuned) 0.724 0.551 0.538 0.584 0.656 0.511 0.591 0.589 0.595 0.417 0.626 0.637
Vanilla (o3-mini) 0.828 0.382 0.632 0.652 0.645 0.542 0.606 0.604 0.366 0.803 0.496 0.491
KATE (o3-mini, 5 shots) 0.884 0.292 0.714 0.708 0.573 0.656 0.555 0.554 0.520 0.474 0.622 0.624
Table 1: Evaluation metrics on the test set for four metrics (F-1, Accuracy, QWK, MSE) across three datasets: MIMIC, KTAS, and ESI-Handbook. Higher values indicate better performance for all metrics except MSE, where lower values are preferable. The best-performing baseline for each metric is bolded, and the second-best and the third-best are underlined. We recommend KTAS as the reference dataset for resolving any discrepancies between findings, due to its large size and gold-standard expert annotations.
Baselines. We implement two ML baselines (Logistic Regression, XGBoost) to see how traditional methods fare in the task of triage. We additionally implement a baseline based on BioBERT (Lee et al. 2020) and a small neural network. Lastly, we implement eight LLM-based approaches: zero-shot with vanilla prompting, zero-shot with CoT prompting (Wei et al. 2022), zero-shot with automatic CoT (AutoCoT) prompting (Zhang et al. 2022), few-shot with vanilla prompting, few-shot with CoT prompting, few-shot using k-nearest neighbors (KNN) retrieval i.e. KATE (Liu et al. 2022), KATE with CoT prompting, and fine-tuning on top of continued pre-training using Llama-3.1-8B-Instruct. Further details on the implementation of our baselines can be found in Appendix A.
@@ 3.2 Counterfactual Analysis of Demographic Bias
Experimental Setup. The counterfactual benchmark is evaluated on prominent LLMs: Llama-3.1-70B-Instruct, Gemini-2.0-Flash, gpt-4o-mini, gpt-4o, claude-3-haiku, claude-3-sonnet, and o3-mini (low reasoning). A temperature of 0 is used for all models. To investigate the impact of reasoning strategies on model bias, we further test most of the models with CoT prompting.
@@
Statistical Analysis. We treat sex and race as independent treatment variables and conduct statistical tests to evaluate the significance of their effects. We apply the Wilcoxon signed-rank test to compare model outputs across male and female variations of the same sample. We apply the Friedman test to evaluate differences in model predictions across the six race categories and their intersections with sex. To account for the testing of multiple LLMs and strategies (e.g. CoT), we apply the Bonferroni correction to control for the false discovery rate in our analysis. For the sake of simplicity, we multiply by 10 rather than 11.
Demographic Llama Gemini gpt-4o-mini gpt-4o claude-3-haiku claude-3-sonnet o3
3.1-70B 2.0 Vanilla CoT Vanilla CoT Vanilla CoT Vanilla CoT mini
Men 2.478 2.860 2.759 2.630 3.018 2.977 2.591 2.894 2.948 3.094 2.869
American Indian 2.458 2.863 2.753 2.666 3.181 2.983 2.800 2.912 3.000 3.132 2.858
Asian 2.468 2.890 2.769 2.612 3.045 3.000 2.500 2.947 3.095 3.066 2.853
Black 2.474 2.895 2.584 2.573 2.931 2.923 2.566 2.883 2.714 3.080 2.879
Hispanic 2.495 2.811 2.846 2.627 2.886 2.940 2.683 2.947 2.928 3.073 2.863
Native Hawaiian 2.458 2.874 2.553 2.565 3.022 3.059 2.216 2.848 2.857 3.161 2.905
White 2.516 2.832 3.046 2.736 3.045 2.957 2.783 2.830 3.095 3.051 2.858
Women 2.486 2.865 2.653 2.629 3.064 2.963 2.436 2.810 2.682 3.068 2.865
American Indian 2.463 2.874 2.692 2.666 3.181 2.889 2.516 2.824 2.785 3.154 2.863
Asian 2.495 2.874 2.707 2.589 3.045 3.042 2.366 2.824 2.642 3.029 2.847
Black 2.479 2.858 2.584 2.627 3.022 2.881 2.333 2.824 2.523 3.022 2.874
Hispanic 2.500 2.868 2.800 2.705 3.022 2.906 2.500 2.824 2.547 3.044 2.890
Native Hawaiian 2.474 2.847 2.384 2.511 3.000 3.093 2.166 2.713 2.595 3.073 2.837
White 2.505 2.868 2.753 2.674 3.113 2.966 2.733 2.853 3.000 3.088 2.879
Sex 1.00 1.00 $<0.01^{**}$ 1.00 1.00 1.00 $<0.01^{**}$ $<0.01^{**}$ $<0.01^{**}$ 1.00 0.728
Race $<0.01^{**}$ 1.00 $<0.01^{**}$ $<0.01^{**}$ $<0.01^{**}$ $<0.01^{**}$ $<0.01^{**}$ 1.00 $<0.01^{**}$ 0.247 0.824
Sex & Race $<0.01^{**}$ 0.39 $<0.01^{**}$ 0.028 0.52 0.01 $<0.01^{**}$ 0.056 $<0.01^{**}$ 0.83 0.693
Table 2: Results of counterfactual analysis, testing prominent LLMs across two configurations: Vanilla Prompting and CoT. The values for each row represent the acuity assigned to patients of that category on average. Bolded values in each column indicate the most prioritized demographic group by the model, while underlined values denote the least prioritized. The row without race represents the averages across race for that particular sex. The final section (the last three rows) provides p-values for statistical significance tests, where * indicates $p<0.05$ and ** indicates $p<0.01$. While the rest have been corrected for multiple testing, we leave the p-values for o3-mini uncorrected to highlight its lack of bias.

### v2

# Joseph Lee1    Tianqi Shang1    Jae Young Baik1    Duy Duong-Tran1 Shu Yang1,*

@@ Abstract
Large Language Models (LLMs) have shown significant promise in clinical decision support, yet their application to triage remains underexplored. In this study, we systematically investigate the capabilities of LLMs in emergency department triage through two key dimensions: (1) robustness to distribution shifts and missing data, and (2) counterfactual analysis of intersectional biases across sex and race. We assess multiple LLM-based approaches, ranging from continued pre-training to in-context learning, as well as conventional machine learning approaches. Our results indicate that LLMs exhibit superior robustness, and we further investigate the key factors contributing to the promising LLM-based approaches. Moreover, in this setting, we also identify critical gaps in LLM preferences that emerge in particular at the intersections of sex and race. LLMs generally exhibit sex-based differences, but they are most pronounced in certain racial groups. These findings suggest that LLMs, while offering powerful capability for clinical triage, encode demographic preferences that may emerge in specific clinical contexts or particular combinations of characteristics, which demand rigorous auditing before real-world integration.
keywords:
Emergency Department Triage; Large Language Models; Artificial Intelligence; Evaluation; Clinical Informatics.
[…]
© 2024 The Authors. Open Access chapter published by World Scientific Publishing Company and distributed under the terms of the Creative Commons Attribution Non-Commercial (CC BY-NC) 4.0 License.
@@ 1 Introduction
Recent advancements of Large language models (LLMs) have shown promise in various clinical settings [1, 2, 3, 4], with example applications ranging from clinical decision support [5] to conversational systems [6]. Amongst these applications, emergency departments (EDs) are some of the most dynamic and high-stakes clinical environments, serving as the initial point of contact for patients requiring urgent medical attention [7]. A critical component of ED operations is the triage process, where patients are prioritized based on clinical urgency to ensure timely and appropriate care. Overtriage misallocates scarce resources and undertriage endangers time-sensitive patients [8]. In an era of historic ED overcrowding, these issues are a significant concern [9]. While triage is an imperfect practice due to limited information [10], triage practices performed by humans are further prone to challenges such as variability in decision-making, lack of training programs, cognitive biases, and decreased quality during overwhelming patient arrivals [11, 8].
@@
Traditional machine learning (ML) algorithms have seen success in clinical triage [12, 13, 14], but they are limited by their reliance on structured input data and sensitivity to distributional shifts. In practice, much of clinical data consists of unstructured text, such as clinical notes, which can be challenging to utilize in conventional ML. Moreover, substantial variation in triage distributions across hospitals has been documented [8], posing additional challenges for model generalisability[3].
In contrast, the pre-trained clinical knowledge [15] of LLMs and their reasoning abilities pose them as a robust model, able to adapt to diverse settings. Beyond mere prediction, LLMs can generate explanatory rationales, serving as a form of second opinion for clinicians and potentially mitigating mispractice [16]. With their potential to improve the efficiency, accuracy, and equity of emergency department triage [17], LLMs are well-positioned to meet the growing demand for AI-driven decision support, an area already seeing early real-world adoption [9, 18].
Despite their potential, the application of LLMs to triage remains an understudied area. Few studies have examined their capabilities [19, 20, 21], and they predominantly focus on basic prompting techniques without testing advanced methodologies. In addition, we find that most studies have not replicated realistic triage scenarios, usually limited to a brief description of symptoms, pain, and vital signs, rather than in-depth clinical notes [9].
Furthermore, a concern comes from the potential biases and inconsistencies in LLMs when applied to triage decisions, where unfair prioritization leads to adverse outcomes against certain demographics. While research has documented various forms of bias in LLMs across different healthcare contexts [22, 23, 24, 25], the implications for hospital triage remain unclear. This knowledge gap is critical given that triage can affect vulnerable patient populations. The high-stakes nature of triage decisions necessitates an extensive assessment of LLMs’ performance across diverse patient populations and clinical scenarios before any potential deployment.
@@
    We comprehensively evaluate across distribution shifts, missing data, and datasets ranging from real hospital records to curated case studies. We assess various LLM-based approaches, ranging from continued pre-training to in-context learning, providing insights for future research directions.
@@ 2 Related Works
LLMs in Biomedical and Clinical Applications. LLMs have been extensively integrated into the biomedical and clinical domain through a range of approaches, from fine-tuning general LLMs with biomedical-specific corpora [26, 27, 28, 29, 30] to equipping frozen LLMs with techniques like chain-of-thought (CoT) reasoning and retrieval-augmented generation (RAG) [31, 32, 33, 34, 35, 36, 37]. In our study, we focus on a few of these recent approaches that have yet to be applied to triage. For example, studies have shown that careful selection of examples can significantly impact model performance [38] and few-shot prompting techniques, such as retrieving k-nearest neighbors (KNN), have been proposed to enhance LLM performance by retrieving relevant examples [39]. Furthermore, recent studies have explored how LLMs acquire knowledge and expertise [40, 41] and we extend these investigations to our specific setting.
@@
LLMs in the Emergency Department. Recent studies have explored LLMs in triage tasks. For example, a study evaluated the performance of LLMs on a binary triage task, comparing two patients and determining which presented higher acuity. However, their approach relied on clinical notes generated after the fact, which would be unavailable at the point of triage [19]. Another study assessed triage accuracy using CoT prompting and found that conditioning on demographic information did not affect model performance. However, the limited scope of their dataset, 45 vignettes, restricts the generalizability of their findings  [42]. Williams extended the study of LLMs to emergency department scenarios such as predicting hospital admission, the need for radiological investigations, and antibiotic prescriptions. The models’ performance remained significantly lower than that of resident physicians across most tasks [43]. The most advanced work to date is an LLM-based agent system that leverages RAG [21]. However, the validity of their findings is constrained by their small sample size and inconsistencies we’ve found in their reported results. In real-world triage, information is limited and frequently incomplete, posing challenges for LLMs operating in dynamic clinical settings. We believe this calls for better benchmarking, which has emerged in other settings [44, 45]. These works represent a crucial step forward, but further research is needed to evaluate the readiness of LLMs.
@@
LLM Bias in Clinical Settings. Recent studies have highlighted significant biases in LLMs, particularly within clinical contexts. For instance, LLMs have been found to exhibit covert rather than overt biases, making such biases less detectable and potentially more harmful[46]. The study identified biases against individuals speaking African American English dialects, illustrating the subtle ways in which LLMs can perpetuate inequities. Another study found that when asking GPT-4 to describe a case of sarcoidosis, the model produced a vignette of a Black patient 97% of the time and a Black female patient 81% of the time [24]. Similarly, Yang employed counterfactual analysis to demonstrate that LLMs’ biased behaviors often align with existing healthcare disparities[25]. For instance, white patients were systematically provided better treatment recommendations than Black patients, despite identical clinical information. However, to date, no research has systematically explored the intersectionality of sex and race in these biases—an important gap that we aim to address in this work.
3 Methods
@@ 3.1 Evaluating the Clinical Utility of LLMs
Datasets. We utilize three datasets: (1) hospital records from MIMIC-IV-ED v2.2 [47], which follow the ESI protocol; (2) hospital records using the Korean Triage Acuity System (KTAS) [48] which were subsequently re-labeled in hindsight by three experts; and (3) curated case studies from the ESI handbook [49].
@@
These datasets differ in two key dimensions. First, these datasets differ in the reliability and nature of their ground truth labels, which affects how we recommend interpreting results based on them. MIMIC labels reflect nurses’ initial assessments, which are prone to mistakes, KTAS labels benefit from expert consensus that was formed after the fact, and the ESI handbook comprises of curated case studies. Second, information availability varies, with KTAS containing additional structured features beyond MIMIC, and the ESI handbook providing narratives with comprehensive details.
Benchmark Creation.We construct our benchmark with carefully controlled train-test distributions. For MIMIC-IV-ED, we partition data by temporal cohorts, using 10,000 patients from 2014–2016 for training and 1,000 patients from 2017–2019 for testing. Furthermore, we apply stratified sampling across missingness (and acuity levels) to ensure representation of incomplete records. For KTAS, we incorporate a domain shift by training on 689 patients from a local hospital and testing on 580 patients from a separate regional hospital, though both are academic, urban medical centers. Finally, for the ESI handbook, we adopt the split from Lu[21].
Baselines. We implement two ML baselines (Logistic Regression, XGBoost) to see how traditional methods fare in the task of triage. Additionally, to incorporate the chief complaint (only available as text), we implement a baseline based on BioBERT [50]. Lastly, we implement eight LLM-based approaches: zero-shot with vanilla prompting, zero-shot with CoT prompting [51], zero-shot with automatic CoT (AutoCoT) prompting [52], few-shot with vanilla prompting, few-shot with CoT prompting, few-shot using k-nearest neighbors (KNN) retrieval i.e. KATE [53], KATE with CoT prompting, and supervised fine-tuning on top of continued pre-training, which is the common recipe to adapt LLMs for a particular domain. Further details on the implementation of our baselines can be found in the later section 3.3.
@@ 3.2 Counterfactual Analysis of Demographic Bias
Experimental Setup. The counterfactual benchmark is evaluated on prominent LLMs: Llama-3.1-70B-Instruct, Gemini-2.0-Flash, gpt-4o-mini, gpt-4o, claude-3-haiku, claude-3-sonnet, and o3-mini (low reasoning). A temperature of 0 is used for all models. To investigate the impact of reasoning strategies on model bias, we further test some of the models with CoT prompting.
@@
Statistical Analysis. We treat sex and race as independent treatment variables and conduct statistical tests to evaluate the significance of their effects. We apply the Wilcoxon signed-rank test to compare model outputs across male and female variations of the same sample. We also apply the Friedman test to evaluate differences in model predictions across the six race categories and their intersections with sex. To account for the testing of multiple LLMs and strategies (e.g. CoT), we apply the Bonferroni correction to control for the false discovery rate in our analysis.
3.3 Implementation Details
Among our baselines, two are notable: KATE, in-context learning with KNN-based retrieval of examples, and fine-tuning on top of continued pretraining. For our implementation, KATE retrieves the $3k$ most similar training examples using BioBERT embeddings of the unstructured text e.g. chief complaints. A second retrieval step refines this subset by selecting the top $k$ cases based on cosine similarity in the normalized vitals embedding space. For fine-tuning on top of continued pretraining, we first extend the unsupervised training of Llama-3.1-8B-Instruct and Qwen3-8B using the ESI handbook for 10 epochs. To prevent data leakage, we exclude all sections that contain training or test questions. This pretraining step is skipped for the KTAS dataset due to differences in triage protocols between ESI and KTAS. Following continued pretraining, we perform supervised fine-tuning on the labelled triage datasets for 25 epochs, except for MIMIC, where we run it for 10 epochs due to its larger dataset size. All fine-tuning is conducted using LoRA [54].
To determine the optimal number of supervised fine-tuning epochs, we evaluated 10, 25, and 100 epochs, finding that 25 epochs provided the best trade-off between performance and overfitting. Across all experiments, for continued pretraining, we use a learning rate of 4e-5, weight decay of 0.01, LoRA rank $r$ = 32, scaling factor $\alpha$ = 32, a cosine learning rate scheduler with a warm-up ratio of 0.1, and a batch size of 8. For task-specific supervised fine-tuning, we adopt a learning rate of 2e-4, weight decay of 0.01, the same LoRA parameters (r = 32, $\alpha=32$), a linear learning rate scheduler with five warm-up steps, and a batch size of 8.
[…]
Evaluation metrics on the test set for four metrics (F-1, Accuracy, QWK, MSE) across three datasets: MIMIC, KTAS, and ESI-Handbook. Higher values indicate better performance for all metrics except MSE, where lower values are preferable. The best-performing baseline for each metric is bolded, and the second-best and the third-best are underlined. We recommend KTAS as the reference dataset for resolving any discrepancies between findings, due to its large size and gold-standard expert annotations. ESI-Handbook KTAS MIMIC Methodology QWK MSE $(\downarrow)$ F-1 Acc. QWK MSE $(\downarrow)$ F-1 Acc. QWK MSE $(\downarrow)$ F-1 Acc. LogReg N.A. N.A. N.A. N.A. 0.369 0.813 0.418 0.472 0.306 0.550 0.483 0.564 XGBoost N.A. N.A. N.A. N.A. 0.059 1.475 0.291 0.342 0.216 1.616 0.211 0.210 BioBERT 0.117 2.904 0.251 0.287 0.074 1.169 0.261 0.330 0.498 0.432 0.640 0.656 Vanilla 0.923 0.213 0.788 0.787 0.599 0.634 0.511 0.520 0.562 0.561 0.585 0.568 AutoCoT 0.932 0.185 0.816 0.815 0.568 0.668 0.493 0.506 0.486 0.665 0.581 0.566 CoT 0.909 0.225 0.809 0.809 0.598 0.660 0.520 0.522 0.386 0.734 0.514 0.510 FewShot (40 shots) 0.940 0.169 0.826 0.831 0.603 0.639 0.516 0.518 0.546 0.494 0.567 0.563 FewShotCoT (40 shots) 0.894 0.292 0.708 0.708 0.624 0.573 0.559 0.561 0.550 0.486 0.591 0.588 KATE (40 shots) 0.934 0.180 0.821 0.820 0.667 0.413 0.696 0.718 0.654 0.352 0.677 0.678 KATECoT (40 shots) 0.917 0.236 0.763 0.764 0.667 0.497 0.607 0.608 0.601 0.402 0.639 0.640 Llama-8B (fine-tuned) 0.724 0.551 0.538 0.584 0.656 0.511 0.591 0.589 0.595 0.417 0.626 0.637 Qwen-8B (fine-tuned) 0.901 0.354 0.693 0.689 N.A. N.A. N.A. N.A. N.A. N.A. N.A. N.A. Vanilla (o3-mini) 0.828 0.382 0.632 0.652 0.645 0.542 0.606 0.604 0.366 0.803 0.496 0.491 KATE (o3-mini, 5 shots) 0.884 0.292 0.714 0.708 0.573 0.656 0.555 0.554 0.520 0.474 0.622 0.624